# Converting sWeights to Probabilities with Density Ratios

### v1

@@ 3.1 Toy Example
Figure 2 shows a comparison of the signal distributions unfolded in $\phi$ and $z$ using the signal sWeights and drWeights for the best algorithm, namely the double density ratio using the GBDT followed by another GBDT. Figure 2 also compares the case where the events were generated with a signal to background ratio of 1:2 or 1:9. A total of 0.5M events were generated events where 0.35M were used in training and 0.15M were used for testing and are plotted in Figure 2. The uncertainties are calculated from the original sWeights. Figure 2 also shows the bin-by-bin pull distributions defined as the difference between the bin values for the two distributions divided by the sum squared of their uncertainties. Overall the density ratio estimation is able to accurately reproduce the $\phi$ and Z signal weighted distributions as shown by the pulls which are largely consistent with being less than 1. This also holds in the case of a dominant background as shown in the bottom row of Figure 2. For both signal to background ratios of 1:2 or 1:9, the $\phi$ asymmetries measured with the signal drWeights were consistent within uncertainties with the generated value of 0.8.
The number of generated events was then varied to ascertain the impact this has on the density ratio estimation using the preferred GBDT and GBDT double density ratio. The results are shown in Table 2, together with the results from the sWeights fitting. The asymmetry amplitude was again generated at 0.8, and here the signal to background ratio was fixed to the large background case at 1:9. In all cases, even with as few as $10^{3}$ generated events, the asymmetry measured with the drWeights is consistent with the generated asymmetry of 0.8. The mean standard deviation and uncertainty on the asymmetry are generally consistent. At 1000 events, corresponding to 100 actual signal events, the fit error is twice the size of the standard deviation, suggesting the uncertainty is overestimated by a factor 2. This is not surprising given that the sum of the squared weights contribution to the uncertainty is less valid at low statistics. Also at 1000 events the drWeights clearly outperform the sWeights. This is due to the latter producing bins with unphysical negative counts, a problem which is resolved by using probability weights. This issue is ultimately an artifact of the weighted binned fitting, which is not the ideal method for extracting parameters from data. Instead one should use an event-by-event maximum likelihood procedure to produce reliable results, but this is outwith the scope of this work. The good performance of the drWeights even with large backgrounds and small event samples is a key consideration, given that experimental datasets may be limited and may have irreducible backgrounds. Although it is worth noting that drWeights should only be used when positive definite probabilities are required as the sWeights proved to be more generally more reliable, overall the drWeights were found to be a robust conversion of sWeights to probabilities.
Number of sWeights sWeights sWeights sWeights drWeights drWeights drWeights drWeights
Events Mean $\sigma$ Uncertainty $\frac{\sigma}{\text{Uncertainty}}$ Mean $\sigma$ Uncertainty $\frac{\sigma}{\text{Uncertainty}}$
$10^{3}$ 17.94 84.81 14.67 5.78 0.679 0.2710 0.5902 0.46
$10^{4}$ 0.870 0.1090 0.0953 1.14 0.778 0.0929 0.1038 0.89
$10^{5}$ 0.804 0.0244 0.0274 0.89 0.793 0.0260 0.0285 0.91
$10^{6}$ 0.799 0.0104 0.0090 1.16 0.792 0.0110 0.0092 1.20
@@
Table 2: Comparison of the $\phi$ asymmetry measured on the sWeights signal distribution and the drWeights (DR) signal distribution when increasing the number of generated events. The signal distribution was generated with a $\phi$ asymmetry of 0.8 and a signal to background ratio of 1 to 9. The data generation and training were repeated 50 times. The mean and standard deviation ($\sigma$) of the measured asymmetries are reported, along with the mean uncertainty. The drWeights are estimated using the GBDT and GBDT double density ratio. Note that to extract the $\phi$ asymmetry, $\phi$ was binned in 100 bins for all tests except at $10^{3}$ events where $\phi$ was binned in 50 bins due to lower statistics.
@@
The toy example presented in this section is contained in the Github repository found at [26]. A generator class produces the total distribution composed of the signal and background distributions, calculates the signal and background sWeights and fits the $\phi$ asymmetry. A plotter class allows to produce the plots shown in this section. A performance class allows to run the tests described here. A trainer class allows to train the models as described in this article. A training script then allows to train the decision tree based models and the neural resampler. The toy example can be used as example of how to train and deploy the methods described in Section 2 and in Ref. [12].
@@ 3.2 $ep\to e^{\prime}\pi^{+}n$ in CLAS12
In this analysis the mass of the missing particles in the reaction $ep\to e^{\prime}\pi^{+}X$ close to the neutron mass was fitted to estimate the number of signal events where the exclusive reaction $ep\to e^{\prime}\pi^{+}n$ was produced. Figure 3 shows the fit of the missing mass of $ep\to e^{\prime}\pi^{+}X$ with the background described by a third order Chebyshev polynomial. The neutron signal was given by a template histogram from simulated data that was created by generating events for the reaction $ep\to e^{\prime}\pi^{+}n$ and running them through the CLAS12 simulation framework, GEMC [30]. This fit allowed sWeights to be assigned to separate the neutron signal from the underlying background. We are now able to disentangle neutron and background distributions. However if we wish to train a machine learning algorithm with these signal neutrons we should convert the sWeights to drWeights as described in Section 2.
@@
Figure 3: An sPlot fit of the invariant mass of $ep\to e^{\prime}\pi^{+}X$ in the region of the neutron missing mass. The background is fitted with a third order Chebyshev polynomial function. The parameters Yld_BG relate to the background yield, with a1, a2 and a3 the Chebyshev polynomial coefficients. The neutron signal is fitted using a simulated template. The parameter Yld_Signal relates to the neutron signal yield. Three parameters, alpha, off, scale, allow to convolute the simulated model to represent the data. The parameter off allows to offset the mean of the distribution, with the scale parameter scaling the width. The parameter alpha allows to convolve the simulated model with an additional Gaussian to better fit the data.
@@
The methodology of Section 2 was applied to the sWeights produced by the fit in Figure 3. We employed a double density ratio with two gradient boosted decision tree (GBDT) steps with a max depth of ten each implemented with the scikit-learn library [19]. The denominator sample was composed of all unweighted events and the numerator sample was composed of all events weighted using the neutron signal sWeights. The GBDTs were trained on the reconstructed neutron spherical momentum components.
Figure 4 shows a comparison of all events, and the same events weighted by the sWeights and the drWeights, for each of the reconstructed momentum magnitude, polar and azimuthal angles. The uncertainties were propagated from the original sWeights. Figure 4 also shows the pull between the two distributions defined as the difference between the bin values for the two distributions divided by the sum squared of their uncertainties. The correlation between momentum and polar angle, and polar and azimuthal angles are also shown. Overall, the distributions are well reproduced when applying the drWeights, as shown by the fact that the pull is consistent with one in nearly all 1 dimensional bins. Accurately converting the sWeights to positive definite probabilities could then have many applications. The drWeights could then be used, for example, to train a machine learning algorithm to model the neutron detection efficiency as a function of all three momentum components reconstructed from the missing four-vector.
@@
Figure 4: Top row: A comparison of all events (black) and all events weighted by the sWeights (blue) or by the drWeights (red) using the density ratio method. The comparison is made for each of the missing momentum (left), polar angle ($\theta$, middle) and azimuthal angle ($\phi$, right). Middle Row: The momentum and the polar angle ($\theta$) correlation histogram weighted using the sWeights (left) or drWeights (middle) and the pull between the two (right). Bottom Row: The polar angle ($\theta$) and azimuthal angle ($\phi$) correlation histogram weighted using the sWeights (left) or drWeights (middle) and the pull between the two (right)
@@ 4 Conclusion And Outlook
Ref. [12] introduced a methodology based on density ratio estimation to resample weights from Monte Carlo event generators so as to avoid negative weights whilst preserving their statistical properties. The key point is that setting up the training sample so as to learn the density ratio of the weights, which in turn allows to resample these, means that the binary cross entropy loss function is not detrimentally affected by negative weights (c.f. Equation 3). This article has detailed how sWeights can be converted to probabilities using density ratio estimation, producing drWeights. We further detailed how gradient boosted decision trees are particularly well suited to converting sWeights. Moreover double density ratios, where a subsequent reweighting is applied to fine-tune the distributions, produces particularly robust weights for the dataset.
@@
Two case studies were presented, one based on a toy example and another on experimental data from the CLAS12 detector. These demonstrated how density ratios can be applied to accurately convert sWeights and preserve their statistical properties. The first case study also highlighted the advantages of using decision trees, namely a non negligible increase in prediction rate and good performance with decreased data sizes and signal to background ratios. The latter is an important consideration as experimental data is limited and may have irreducible signal to background ratios. It is generally beneficial to have several options of learning algorithms to choose from depending on the dataset and task at hand.
@@
The methodology presented in this article can be used for many applications of machine learning to experimental data in high energy and nuclear physics. In general, it is necessary to disentangle the distributions of different event species in experimental data to be able to use the distributions of specific species to train a machine learning algorithm. The second case study presented here detailed how converting sWeights to drWeights could be applied as a useful first step to learning the neutron detection efficiency in the CLAS12 forward detector. The mass of missing particles in $ep\to e^{\prime}\pi^{+}X$ can be fitted to distinguish the neutron signal from the background. The signal sWeights can then be converted to drWeights as described in Section 3.2. These converted weights can then be used in training a binary classification algorithm to model the CLAS12 forward detector neutron detection efficiency. Previous work has demonstrated the use of density ratio estimation via binary classification to model detector efficiency from simulation [10, 11]. The method to obtain drWeights described in this article could be a first step to modelling detector efficiency from experimental data. In general, many such applications will exist and require to accurately convert sWeights to positive definite probabilities as done here.
@@
Finally we note that for general analysis, for example extracting an asymmetry from a background subtracted distribution, sWeights themselves will still provide a more reliable method. The goal here is not to replace sWeights in these circumstances, only in the cases such as machine learning training where positive definite probabilities are required.

### v2

@@ 3.1 Toy Example
Figure 2 shows a comparison of the signal distributions separated from background in $\phi$ and $z$ using the signal sWeights and drWeights for the best algorithm, namely the density ratio product using the GBDT followed by another GBDT. Figure 2 also compares the case where the events were generated with a signal to background ratio of 1:2 or 1:9. A total of 0.5M events were generated events where 0.35M were used in training and 0.15M were used for testing and are plotted in Figure 2. The uncertainties are calculated from the original sWeights. Figure 2 also shows the bin-by-bin pull distributions defined as the difference between the bin values for the two distributions divided by the sum squared of their uncertainties. As the uncertainties are equal for both histograms, given that they come from the sum of the sWeights squared, the pull calculation simplifies to:
$g_{hist}=\frac{S_{i}-DR_{i}}{\sqrt{2}\sigma_{S}}.$ (9)
[…]
In order to study the statistical properties of the drWeighted distributions in more detail, we ran a longer series of 2000 toy test iterations, each with $10^{5}$ generated events and a signal to background ratio of 1:9. The results of this are shown in Table 1. We see that the sWeights perform as expected with the fits producing an unbiased mean with an average fit uncertainty, $\bar{\sigma}_{fit}$, equal to the standard deviation, $\hat{\sigma}_{rms}$, of the $10^{5}$ results. On the other hand, we see that there is a small bias in the drWeighted distribution, with its mean amplitude reduced by $1\%$ and $\hat{\sigma}_{rms}$ increasing moderately by $15\%$. The bias is around one-quarter of a standard deviation. Further we quote the mean reduced $\chi^{2}$ of the 2000 fits; again we see that the sWeighted distributions have values close to 1 as expected, while the drWeighted distributions are significantly lower at 0.66. This is a result of bin-by-bin smoothing resulting from the converter, which must effectively average over neighboring events to assess the density ratio. This can be seen in Figure 2 where the drWeights distributions have less statistical fluctuations. We also show the average of the pull means and widths, where the individual pull means and widths were calculated from the weighted histogram content ($H_{i}$) and fit result ($f_{i}$) :
$g_{fit}=\frac{H_{i}-f(\phi_{i})}{\sigma_{i}}.$ (10)
The mean and standard deviation of $g_{fit}$ are then averaged over all 2000 fits. The sWeights distributions follow a normal distribution, while the drWeights distributions have a significantly lower width of 0.81 (compared to unity), again a result of the smoothing introduced by the density ratio converter.
Weights
Amp. Mean $\hat{\sigma}_{rms}$ $\bar{\sigma}_{fit}$ $\frac{\hat{\sigma}_{rms}}{\bar{\sigma}_{fit}}$ $\chi^{2}$ $\bar{g_{fit}}$ $\bar{\sigma}_{g_{fit}}$
sWeights
0.801 0.027 0.027 1.01 0.94 0.02 0.97
drWeights
0.792 0.031 0.027 1.15 0.66 0.01 0.81
@@
Table 1: Comparison of the $\phi$ amplitude measured with the signal distribution generated with a $\phi$ amplitude of 0.8 for $10^{5}$ events with a signal to background ratio of 1:9. The data generation and training were repeated 2000 times. The mean and standard deviation ($\hat{\sigma}_{rms}$) of the measured amplitudes are reported, along with the mean fit uncertainty. We also show the mean $\chi^{2}$ of the 2000 fits and the average of the pull ($g_{fit}$) means and standard deviations of the histogram to the fit result, with each mean and standard deviation constructed from the 100 bins of the test histogram.
To further investigate the effect of converter smoothing, we construct a covariance matrix for the two types of weighted distribution. This is done by saving the individual histogram bin contents for each test, such that we have a dataset of 100 by 2000 entries. The numpy.cov routine [26] was then used to produce a 100x100 matrix from these data. The covariance matrices for sWeights and drWeights are shown in Figure 3. Note that, for visualisation purposes, the colour axis was capped to $\pm 300$ whereas the actual maximal covariance was of order 800. The sWeights matrix is shown to be diagonal as expected for histograms with uncorrelated bins. Bins around the peaks and troughs of the distribution in the drWeights matrix have a significant correlation with their nearest neighbors which also extends to more distant neighbors. This smoothing of the peaks and troughs is then what leads to the small bias in the fit amplitude. These correlations will also result in $\hat{\sigma}_{rms}$ being larger than $\bar{\sigma}_{fit}$, as the full covariance matrix is not used in the $\chi^{2}$ calculation. In principle, one could correct for that using the full covariance matrix, although that would need to be calculated numerically similar to what is done here.
Figure 3: The sWeight and drWeights bin-to-bin covariance matrices in $\phi$ estimated over 2000 test iterations. The colour axis is capped between -300 and 300 for visualisation purposes but the actual maximal covariance was of order 800.
We then used the bootstrap technique in order to reliably extract the uncertainty in the drWeighted fits for a single data set: for the chosen data set of $10^{5}$ events, the same number of events were randomly sampled from that original set. The sampling is done with replacement, meaning that the same event may appear multiple times or never. This random sampling was performed 2000 times each on two separate data sets and the results are shown in Table 2. The expectation is that sWeights and drWeights reproduce the $\phi$ amplitude in the chosen data set in a given bootstrap sample, following a statistical distribution of 0.8 $\pm\hat{\sigma}_{rms}$, where $\hat{\sigma}_{rms}$ was estimated with independent toy samples and that $\hat{\sigma}_{boot}=\hat{\sigma}_{rms}$. We observe that the mean amplitude is within around one standard deviation of the true value. As before, the $\hat{\sigma}_{boot}$ of the drWeights increases relative to the $\bar{\sigma}_{fit}$. The sWeights mean fit reduced $\chi^{2}$ of the 2000 samples is increased beyond unity as one would expect, due to the bootstrap sampling in each set of events. This variation is also reflected in the pull widths, $\bar{\sigma}_{g_{fit}}$ increasing beyond 1. We also see that the impact of the bin-by-bin smoothing introduced by drWeights is similar to that on independent toy samples. The mean fit reduced $\chi^{2}$ and the pull widths are closer to 1 for drWeights as the increases seen in sWeights are canceled to some degree by the smoothing in the density ratio. Overall we demonstrate that drWeights are well suited to real world applications where precise statistical evaluations are not necessary.
Weights
Amp. Mean $\hat{\sigma}_{boot}$ $\bar{\sigma}_{fit}$ $\frac{\hat{\sigma}_{rms}}{\bar{\sigma}_{fit}}$ $\chi^{2}$ $\bar{g_{fit}}$ $\bar{\sigma}_{g_{fit}}$
sWeights
0.798 0.0270 0.0269 1.00 1.88 0.03 1.38
0.837 0.0282 0.0278 1.01 1.78 0.04 1.34
drWeights
0.784 0.0322 0.0269 1.19 1.12 0.02 1.06
0.806 0.0320 0.0292 1.10 1.03 0.02 1.02
Table 2: Comparison of the $\phi$ amplitude measured on 2000 bootstrap samples drawn from data sets with the signal distributions generated with a $\phi$ amplitude of 0.8 for $10^{5}$ events and signal to background ratios of 1:9. Note that the bootstrapping was done on two different data sets corresponding to the two rows for both sWeights and drWeights. The mean and standard deviation ($\hat{\sigma}_{boot}$) of the measured amplitudes are reported, along with the mean fit uncertainties. We also show the mean $\chi^{2}$ of the 2000 fits and the average of the pull ($g_{fit}$) means and standard deviations, with each mean and standard deviation constructed from the 100 bins of the test histogram.
@@
The toy example presented in this section is contained in the Github repository found at [27]. A generator class produces the total distribution composed of the signal and background distributions, calculates the signal and background sWeights and fits the $\cos(2\phi)$ amplitude. A plotter class allows to produce the plots shown in this section. A performance class allows to run the tests described here. A trainer class allows to train the models as described in this article. A training script then allows to train the decision tree based models and the neural resampler. The toy example can be used as example of how to train and deploy the methods described in Section 2 and in Ref. [12].
3.2 Higher Frequency Distributions
In Section 3 we showed that for a relatively common and simple problem in physics data analysis the drWeights were able to reproduce signal distributions with reasonably well behaved statistical properties. However, given that the underlying GBDT algorithm effectively relies on averaging over local densities, some degree of smoothing, or correlating of events, is to be expected, as was demonstrated at the end of Section 3. To investigate the limits of the drWeights technique, we also performed a toy analysis where the structures in the distributions are sharper and so smoothing effects are likely to be more noticeable. This was done by gradually increasing the frequency of the cosine distribution and performing the same toy analysis.
In Figure 4 we show the full event, sWeight and drWeight distributions for a cosine of frequency 10, for the two different signal-to-background ratios we used previously. Visually, the sWeight and drWeight distributions agree well and the pull distributions are still within expected limits. On closer inspection, it is evident that the larger background sample has slightly larger peaks and dips for the blue sWeight distribution than the red drWeight one. This leads to the drWeight events having a smaller fitted amplitude than for the sWeights.
Figure 4: A comparison of all events, sWeights signal events (blue) and drWeights red for a cosine of frequency = 10, a for 1:2 (left) and 1:9 (right) signal-to-background ratio.
In Figure 5 we systematically increase the frequency from 2 to 10 and report the fitted amplitude averaged over 50 fits, for the two types of weights and two signal-to-background ratios. We observe a steady fall in the drWeights amplitudes that is significantly stronger for the larger background sample. This highlights the effect of the smoothing in the algorithm; it has a similar effect to adding experimental smearing. It is therefore important when deciding to use this method to evaluate if such limitations will significantly affect the results in a particular use-case.
Figure 5: Effect on the fitted cosine amplitude from increasing its frequency. We compare the sWeight and drWeight cases for different signal-to-background ratios.
As shown in table 4, the density ratio product performs consistently better than a single density ratio. This is due to the second reweighting step correcting the bias introduced by the smoothing of the first model. We then increased the number of converters from 2 to 25. Figure 6 shows the impact on the drWeights amplitudes, along with a comparison between the uncertainty and the standard deviation, for a cosine frequency of 10. Adding further converters allows to correct the amplitude up until the drWeights reproduce the sWeights amplitude. However the standard deviation becomes larger than the uncertainty, pointing to the fact that there is larger than expected fluctuation in results between training iterations. From a practical perspective this should not be an issue when using the drWeights method, for example to create datasets from experimental for the purposes of training machine learning algorithms. The training of the drWeights converters can be repeated many times with the best iteration used in further applications. Training fewer density ratio converters will produce stabler results, and decreases the training and prediction times. A single density ratio product should therefore be preferred over many density ratios, but this may be a viable solution in complicated scenarios as shown here.
Figure 6: Effect on the fitted cosine amplitude from increasing the number of GBDTs, for a cosine frequency of 10. We compare the sWeight and drWeight cases and the uncertainty and standard deviations of the drWeights.
The improvement in the drWeights amplitude is due to the subsequent reweighting steps correcting the smearing introduced by previous steps which smoothed over features in the signal distribution. However, it must be understood that adding many more reweighters injects further correlations between events through the estimation of the local density ratio, to better reproduce the sWeighted distributions and reduce bias. In noisy data, this procedure can lead to more stable results, as we effectively smooth out the noise if the induced bias is less than the noise effects. An example of this is shown in Table 5 in Appendix A, for the low event-number tests. In contrast, sWeights are model agnostic and simply provide the best unbiased estimate of the control variables given the a priori knowledge available for the discriminating variables. The drWeights are therefore a useful tool to reproduce the sWeighted distributions, but they are by construction biased.
The fact that the standard deviation becomes larger than the uncertainty when using many drWeights converters also points to some limitations in the statistical properties of drWeights, especially in scenarios such as the one described above with sharp features and large backgrounds.Table 4 showed that the uncertainty in the data sets produced with drWeights, which therefore have correlated bins due to the smoothing and bias described above, is generally somewhat larger than that obtained from an uncorrelated dataset, for example with sWeights. In Section 2.2 we suggested carrying over the uncertainty from the sWeights uncertainty calculation; our results show that this may, however, underestimate the uncertainty on the drWeights distributions, perhaps due to a lack of precision inherent in the training procedure. Therefore, in general, sWeights should be preferred in cases where their well defined statistical properties and uncertainties are required. In turn, drWeights should be used for practical applications where positive definite probabilities are required, such as creating machine learning training samples from experimental data containing contributions from several different event sources.
@@ 3.3 $ep\to e^{\prime}\pi^{+}n$ in CLAS12
Figure 7: An sPlot fit of the invariant mass of $ep\to e^{\prime}\pi^{+}X$ in the region of the neutron missing mass. The background is fitted with a third order polynomial function. The parameters Yld_BG relate to the background yield, with a1, a2 and a3 the polynomial coefficients. The neutron signal is fitted using a simulated template. The parameter Yld_Signal relates to the neutron signal yield. Three parameters, alpha, off, scale, allow to convolute the simulated model to represent the data. The parameter off allows to offset the mean of the distribution, with the scale parameter scaling the width. The parameter alpha allows to convolve the simulated model with an additional Gaussian to better fit the data.
In this analysis the mass of the missing particles in the reaction $ep\to e^{\prime}\pi^{+}X$ close to the neutron mass was fitted to estimate the number of signal events where the exclusive reaction $ep\to e^{\prime}\pi^{+}n$ was produced. Figure 7 shows the fit of the missing mass of $ep\to e^{\prime}\pi^{+}X$ with the background described by a third order polynomial. The neutron signal was given by a template histogram from simulated data that was created by generating events for the reaction $ep\to e^{\prime}\pi^{+}n$ and running them through the CLAS12 simulation framework, GEMC [31]. This fit allowed sWeights to be assigned to separate the neutron signal from the underlying background. We are now able to disentangle neutron and background distributions. However if we wish to train a machine learning algorithm with these signal neutrons we should convert the sWeights to drWeights as described in Section 2, by training a density ratio converter with their three momentum components as input features.
@@
The methodology of Section 2 was applied to the sWeights produced by the fit in Figure 7. We employed a density ratio product with two gradient boosted decision tree (GBDT) steps with a max depth of ten each implemented with the scikit-learn library [19]. The denominator sample was composed of all unweighted events and the numerator sample was composed of all events weighted using the neutron signal sWeights. The GBDTs were trained on the reconstructed neutron spherical momentum components.
@@
Figure 8: A comparison of all events (black) and all events weighted by the sWeights (blue) or by the drWeights (red) using the density ratio method. The comparison is made for each of the missing momentum (left), polar angle ($\theta$, right) and azimuthal angle ($\phi$, bottom).
Figures 8 and 9 show a comparison of all events, and the same events weighted by the sWeights and the drWeights, for each of the reconstructed momentum magnitude, polar and azimuthal angles. The uncertainties were propagated from the original sWeights. These figures also show the pull between the two distributions defined as the difference between the bin values for the two distributions divided by the sum squared of their uncertainties. The correlation between momentum and polar angle, and polar and azimuthal angles are also shown. Overall, the distributions are well reproduced when applying the drWeights. Accurately converting the sWeights to positive definite probabilities could then have many applications. The drWeights could be used, for example, to train a machine learning algorithm to model the neutron detection efficiency as a function of all three momentum components reconstructed from the missing four-vector.
Figure 9: Top Row: The momentum and the polar angle ($\theta$) correlation histogram weighted using the sWeights (left) or drWeights (middle) and the pull between the two (right). Bottom Row: The polar angle ($\theta$) and azimuthal angle ($\phi$) correlation histogram weighted using the sWeights (left) or drWeights (middle) and the pull between the two (right).
@@ 4 Conclusion And Outlook
Ref. [12] introduced a methodology based on density ratio estimation to resample weights from Monte Carlo event generators so as to avoid negative weights. The key point is that setting up the training sample so as to learn the density ratio of the weights, which in turn allows to resample these, means that the binary cross entropy loss function is not detrimentally affected by negative weights (c.f. Equation 3). This article has detailed how sWeights can be converted to probabilities using density ratio estimation, producing drWeights. We further detailed how gradient boosted decision trees are particularly well suited to converting sWeights. Moreover density ratio products, where a subsequent reweighting is applied to fine-tune the distributions, produces particularly robust weights for the dataset.
@@
Three case studies were presented, one based on a simple toy example, the second extending this to more complicated situations and another on experimental data from the CLAS12 detector. These demonstrated the potential of density ratios to convert sWeights. The first case study also highlighted the advantages of using decision trees, namely a non negligible increase in prediction rate and good performance with decreased data sizes and signal to background ratios. The latter is an important consideration as experimental data is limited and may have irreducible signal to background ratios. It is generally beneficial to have several options of learning algorithms to choose from depending on the dataset and task at hand.
However, the first case study also demonstrated that the drWeights converter introduces some degree of smoothing, or correlating of events, as the converter averages over local densities when producing the drWeights. The second case study showed that when backgrounds are large and distribution features are sharp, these smoothing effects can lead to significant loss in fidelity. This can be compensated by introducing many subsequent reweighting steps at a loss in training stability and training and prediction times. As such, care should be taken in evaluating the applicability and implementation of the density ratio weight conversion to given use cases. Nevertheless, the drWeights are generally a useful tool for converting sWeights to probabilities.
@@
The final case study demonstrated how the density ratio conversion of sWeights can be well suited to real experimental data. The methodology presented in this article can be used for many applications of machine learning to experimental data in high energy and nuclear physics. In general, it is necessary to disentangle the distributions of different event species in experimental data to be able to use the distributions of specific species to train a machine learning algorithm. For example, converting sWeights to drWeights as shown in the third case study could be applied as a useful first step to learning the neutron detection efficiency in the CLAS12 forward detector. The mass of missing particles in $ep\to e^{\prime}\pi^{+}X$ can be fitted to distinguish the neutron signal from the background. The signal sWeights can then be converted to drWeights as described in Section 3.3. These converted weights can then be used in training a binary classification algorithm to model the CLAS12 forward detector neutron detection efficiency. Previous work has demonstrated the use of density ratio estimation via binary classification to model detector efficiency from simulation [10, 11]. The method to obtain drWeights described in this article could be a first step to modeling detector efficiency from experimental data. In general, many such applications will exist and require to accurately convert sWeights to positive definite probabilities as done here.
@@
Finally we note that for general analysis, for example extracting an amplitude from a background subtracted distribution, sWeights themselves will still provide a more reliable method. The goal here is not to replace sWeights in these circumstances, only in the cases such as machine learning training where positive definite probabilities are required.